OpenStudy (anonymous):

solve the identity: sin(x-3pi/2)=cosx

OpenStudy (invinceble):

idk

OpenStudy (gorv):

open the sin

OpenStudy (gorv):

sin(A-B)=sinA*cosB-sinB*cosA

OpenStudy (gorv):

A=x b=3*pi/2

OpenStudy (gorv):

plug in values of A and B

OpenStudy (anonymous):

sin(x)cos(0)-sin(0)0cos(x)

OpenStudy (anonymous):

sorry, there's an extra 0

OpenStudy (gorv):

B=3pi/2 not zero

OpenStudy (gorv):

\[sinx*\cos \frac{ 3 *\pi}{ 2 }-\sin \frac{ 3*\pi }{ 2 }*cosx\]

OpenStudy (gorv):

what is cos value at cos(3pi/2)=???